OpenStudy (anonymous):

help? If y varies inversely with x and y = 8 when x = 5, then what is y when x = 2?

OpenStudy (anonymous):

y= c/x.

OpenStudy (anonymous):

8=c/5, so c=40.

OpenStudy (anonymous):

y=40/x, so y=40/2=20 when x=2.